OpenStudy (anonymous):

No clue how to solve can somoene please explain. Let f and g be differentiable functions with following properties 1)g(x)>0 for all x 2)f(0)=1 If h(x)=F(x)g(x) and h'(x)=f(x)g'(x), then f(x)= a)f'(x) b)g(x) c)e^x d)0 e)1

OpenStudy (anonymous):

if h(x)=f(x)g(x) then generally h'(x)=f'(x)g(x)+f(x)g'(x), so in this case i think the answer would be e: f(x)=1

OpenStudy (anonymous):

ooh so you use product rule?

OpenStudy (anonymous):

right, e^x looks close but its not the product rule if f(x)=e^x then with h(x)=e^xg(x) then h'(x)=e^xg(x) + e^xg'(x)

OpenStudy (anonymous):

if f(x) = 1 then h'(x) = 1* g'(x)+0g(x)

OpenStudy (anonymous):

ah get it now thanks soo much

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Nice!